The mean height of the grade nine students at a certain high school is 164 centimeters and the

standard deviation is 10 centimeters. Assuming the heights are normally distributed, what

percent of the heights is greater than 168 centimeters?


1
Expert's answer
2022-04-26T00:51:47-0400

We have a normal distribution, "\mu=164, \sigma=10."

Let's convert it to the standard normal distribution,

"z=\cfrac{x-\mu}{\sigma}=\cfrac{168-164}{10}=0.40;"


"P(X>168)=P(Z>0.40)=\\
=1-P(Z<0.40)=\\
=1-0.6554=0.3446=34.46\%\text{ (from z-table)}."